OpenStudy (anonymous):

What is the equation of the line passing through (2, –2) and perpendicular to the line y = –x + 3 in slope-intercept form?

OpenStudy (ajprincess):

Product of slopes of perpendicular lines=-1 \(y-y_1)=m(x-x_1)\)

OpenStudy (anonymous):

y=x-2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome. Do you understand how i got that?